OpenStudy (nbarrera):

Help Please+MEDAL! The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 13, c = 21 A. 15.87 B. 16.19 C. 15.83 D. 16.49

OpenStudy (mathstudent55):

In a right triangle with legs of lengths a and b, and hypotenuse of length c, the Pythagorean theorem states that \(a^2 + b^2 = c^2\) You know b and c. Plug them in the Pythagorean equation above, and solve for a.

OpenStudy (anonymous):

Which is a^2 + b^2 = c^2

OpenStudy (anonymous):

its D

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (nbarrera):

thx:)